OpenStudy (anonymous):

Fan and Medal Kenny mowed a lawn today to earn some extra money. He spent $1.40 on fuel and $3.25 on lunch. He charged $15 to mow the lawn. What was his profit for mowing the lawn?

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

you add the cost of lunch and gas together which is 4.65 and than subtract than from 15 which is 10.35

OpenStudy (anonymous):

Thank you can you please help me with a few more?

OpenStudy (anonymous):

ill try

OpenStudy (anonymous):

Nolan lost 12.5 points in the final round of a game, bringing his score down to –8. Suppose p represents Nolan’s points before he lost the final round. Which equation can be used to represent this situation? A. p – (–8) = –12.5 B. p + 12.5 = –8 C. p – 12.5 = –8 D. p + 8 = 12.5

OpenStudy (anonymous):

its d

OpenStudy (anonymous):

A. p -(-8)= _12.5 B. p + 12.5 = –8 C. p – 12.5 = –8 D. p + 8 = 12.5

OpenStudy (anonymous):

one more

OpenStudy (anonymous):

On Wednesday, William wrote a check in the amount of $45.60. This brought his balance down to –$15.75. Suppose n represents William’s balance before he wrote the check. Which equation can be used to represent this situation? A. n + 15.75 = –45.60 B. n – 15.75 = 45.60 C. n + 45.60 = –15.75 D. n – 45.60 = –15.75

OpenStudy (anonymous):

is d n-45.60= 15.75 or is it n45.60=15.75?

OpenStudy (anonymous):

n-45=-15.75

OpenStudy (anonymous):

D.

OpenStudy (anonymous):

thn its d

OpenStudy (anonymous):

the one before this is c i hit the wrong bottom

OpenStudy (anonymous):

I got a 60

OpenStudy (anonymous):

what score do you have to get?

OpenStudy (anonymous):

80 or higher

OpenStudy (anonymous):

hmm. i wonder why its a 60. maybe the @grray7 could lookat all the question tell us if i was wrong?

OpenStudy (anonymous):

i missed the questions that i answered

OpenStudy (anonymous):

Solve the word problem. The table shows the low outside temperatures for Monday, Tuesday, and Wednesday. What was the change in the low outside temperature from Tuesday to Wednesday? This function table shows the outside temperature for a given day A. –20.8⁰ B. –12⁰ C. –4.2⁰ D. –3.2⁰

OpenStudy (anonymous):

tuesday to wednesday is -8.8 -12

OpenStudy (anonymous):

a. -20.8 b.-12 c.-4.2 d. -12

OpenStudy (anonymous):

you still there

OpenStudy (anonymous):

its a

OpenStudy (anonymous):

On Monday, Lucy’s bank balance was –$164. On Tuesday, her bank balance was –$17. What was the change in Lucy’s bank balance from Monday to Tuesday?

OpenStudy (anonymous):

they are negative

OpenStudy (anonymous):

both of them or just one of them?

OpenStudy (anonymous):

both

OpenStudy (anonymous):

she paid 147 dollars

OpenStudy (anonymous):

three more

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

Sarah earned some money taking care of a dog. She charged $8.50 for washing the dog and $7 for walking the dog. She spent $1.15 on shampoo to wash the dog. What was Sarah’s profit from taking care of the dog? sorry 2 more

OpenStudy (anonymous):

$14.35

OpenStudy (anonymous):

yes that right

OpenStudy (anonymous):

thanks yall

OpenStudy (anonymous):

your welcome

OpenStudy (anonymous):

Need help with anything else?

OpenStudy (anonymous):

The temperature decreased overnight by 15.2°C to a temperature of –12.8°C. Suppose t represents the temperature before it decreased. Which equation can be used to represent this situation? A. t – 15.2 = –12.8 B. t + 12.8 = –15.2 C. t – 12.8 = 15.2 D. t – (–15.2) = –12.8

OpenStudy (anonymous):

one more after this

OpenStudy (anonymous):

i dont know that one

OpenStudy (anonymous):

Me neither. Sorry.

OpenStudy (anonymous):

a.12.8 is negative b.15.2 is negative d. 12.8 is negative d. 15.2 is negative

OpenStudy (anonymous):

i have to get off the computer.

OpenStudy (anonymous):

ok one more

OpenStudy (anonymous):

A zoologist is recording the loss of wolves in a her state. She notes the number of wolves, w, in the state on January 1. One year later, there were 84 wolves in the state, which is 23 fewer wolves than were in the state a year earlier. Which equation represents this situation? A. w + 23 = 84 B. w – 23 = 84 C. 23 – w = 84 D. w – 84 = negative 23

OpenStudy (anonymous):

b is correct I took the assessment